(a)Derive an expression for VA − VB, the potential difference between points A and B (b) Does your result reduce to the expected answer when d = 0? When a = 0? and when q = 0.


1
Expert's answer
2021-07-04T17:45:39-0400

"V_A=\frac{kQ}{r_A},V_B=\frac{kQ}{r_B}"


"V_A-V_B=\frac{kQ}{r_A}-\frac{kQ}{r_B}={kQ}({\frac{1}{r_A}-\frac{1}{r_B}})"

"V_A-V_B=kQ\frac{r_B-r_A}{r_Ar_B}"

When

"d=r_B-r_A=0;q=Q=0;a=r_A=0"

Put value then We can written as

"V_B-V_A=0"

"V_A=V_B"